A semicircular window of radius 16 inchesis to be laminated with a sunblock coating that costs $0.65 per square inch to apply. What is the total cost of coating the window, to the nearest cent? (Use the Area of a Circle and Pi = 3.14)
Answer by stanbon(75887)

A semicircular window of radius 16 inches is to be laminated with a sunblock coating that costs $0.65 per square inch to apply. What is the total cost of coating the window, to the nearest cent? (Use the Area of a Circle and Pi = 3.14)
--------------
Area = pi*r^2
Semi-Area = (1/2)pi*r^2
SA = (1/2)(3.14)*8^2 = 100.48 sq.in.
------------------------------------
Cost = 0.65*100.48 = $65.31
==================================
Cheers,
Stan H.